Calculate the cell potential for the following reaction:


Ni(s) + 2 Cu+(0.010M) → 2 Cu(s) + Ni+2(0.0010M)


  1. 0.40 V
  2. -0.43 V
  3. 0.43 V
  4. 0.34 V
  5. 0.77 V
1
Expert's answer
2021-09-19T00:03:21-0400

At anode (Oxidation reaction occur)

Ni (s)-------->Ni2+(aq) + 2e-

E0cell= -0.25 V

At cathode Reduction reaction occur

2Cu+(aq) + 2e- ------->2Cu(s).

E0cell= 0.520 V

So overall standard cell potential is

Eocell= Eocell reduction - Eocell oxidation

Eocell = 0.520-(-0.25)

=0.77V


Ecell =Eocell - 0.059/n{log[Ni2+]/[Cu+]2}

n= no.of change of electron

Ecell=0.77-0.059/2{log[0.0010]/[0.0102]}

Ecell = 0.77-(0.0591/2)log(10)

Ecell=0.77-(0.0591/2)(1)

Ecell=0.77-0.0295

Ecell=0.7405V